OpenStudy (anonymous):

can someone please show me step by step how to get the point slope and slope intercept form for x-intercept= -2/3 and y-intercept= 6? cause i kepp gettingy-3=2/3 (x+2) for point slope and y=2/3x+11/3

OpenStudy (anonymous):

start by finding the slope with the intercepts (0, -2/3) and (6,0)

OpenStudy (anonymous):

i did that

OpenStudy (anonymous):

but some how i still mess up

OpenStudy (anonymous):

the equation for slope should look like this \[\frac{ y2-y1 }{ x2-x1 } = \frac{ 0--\frac{ 2 }{ 3 } }{ 6-0 } = \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

oh i see what i did i had the wrong one on top, duh fail lol. Thanks